OpenStudy (anonymous):

The students at Midtown Middle school sold flowers as a fundraiser in September and October. In October, they charged $1.50 for each flower. The October price was a 20% increase of the September price. Part A What was the price of the flowers in September? Enter your answer in the box. $ Part B The seventh-grade class earned 40% of the selling price of each flower. In September, they sold 900 flowers. In October, they sold 700 flowers. Select a choice from each drop-down menu to make a true statement. The seventh-grade class earned more money in by

OpenStudy (anonymous):

ill give u a mete and fan

OpenStudy (krazykricket):

do you know percentage already?

OpenStudy (anonymous):

no

OpenStudy (krazykricket):

ok when we say 20%, it means 20 per 100 or 20/100 or 0.2...

OpenStudy (anonymous):

7th grader

OpenStudy (krazykricket):

ok for Part A... if P is the price for September, $1.50 for October, it is the 20% increase of P in September, therefore we solve for P, the equation which I want you to solve is P+0.2P=1.50

OpenStudy (krazykricket):

all you have to do is perform the arithmetic operations on like terms such as P, what is 1 + 0.2?

OpenStudy (anonymous):

1.2

OpenStudy (krazykricket):

so you'll have1.2\times P=1.50now what is the reverse of multiplication, that is the operation for you to solve P... you get 1.50 when you 1.2 to a number P... so how will you solve P?

OpenStudy (anonymous):

1.2xp=1.5

OpenStudy (krazykricket):

solve for p

OpenStudy (anonymous):

0.8

OpenStudy (krazykricket):

it seems not... how did you perform it...

OpenStudy (krazykricket):

P=\frac{1.5}{1.2}

OpenStudy (anonymous):

division

OpenStudy (krazykricket):

so P = $1.25 for September, 0.25 is the 20% of it... so if we increase 20% by October, it will be 1.25 + 0.25 = 1.50... hope you understand it...

OpenStudy (anonymous):

so the answer is 0.8?

OpenStudy (anonymous):

yes or no?

OpenStudy (krazykricket):

Yes

OpenStudy (krazykricket):

1.50